OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Dolphin Tricks & Tours. Age Ticket Price Under three years free Three to twelve years $35 Over twelve years $46 On a certain day, the company took 88 people on whale watching trips. There were 44 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

(44 + x)35 + (88 – (44 + x)) · 46 = C (88 - x)35 + (88 – 44) · 46 = C (88 + x)35 + (88 – (44 + x)) · 46 = C (44 – x)35 + (88 – 44) · 46 = C

OpenStudy (anonymous):

those are the possivble 4 answers it could be

OpenStudy (anonymous):

what do you think the answer is ?

OpenStudy (anonymous):

She put down the last one this is her homework not mine.

OpenStudy (anonymous):

Yeah i think that's right

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

Its worth 8 points off of her homework

OpenStudy (anonymous):

is she online school

OpenStudy (anonymous):

Yeah. She refuses to go to a regular school.

OpenStudy (anonymous):

Nevermind its C and lol i do flvs to

OpenStudy (anonymous):

Okay wanna know a secret.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I dont have a lil sis its me. Im good at it Im just checking my work. Your the first kid Ive actually met in flvs lol you should add me on facebook I have like no friends.

OpenStudy (anonymous):

haha true and yeah whats your email

OpenStudy (anonymous):

JoeBlowCat@yahoo.com

OpenStudy (anonymous):

i sent it

OpenStudy (anonymous):

I accepted it. XD lol yeah so hmm idk what to say

OpenStudy (anonymous):

lol what garde you in ?

OpenStudy (anonymous):

Well lol look what I said on facebook.

OpenStudy (anonymous):

So what's the answer?? XD

OpenStudy (anonymous):

c

OpenStudy (anonymoushd):

welp :/